# BotArtist: Generic approach for bot detection in Twitter via semi-automatic machine learning pipeline

### v3

@@ Abstract
Twitter, as one of the most popular social networks, provides a platform for communication and online discourse. Unfortunately, it has also become a target for bots and fake accounts, resulting in the spread of false information and manipulation. In this paper, we introduce a semi-automatic machine learning pipeline (SAMLP) designed to address the challenge of detecting Twitter bots. Through this pipeline, we have developed a comprehensive bot detection model named BotArtist, based on user profile features. SAMLP leverages nine distinct publicly available datasets to train the BotArtist model. To assess BotArtist’s performance against current state-of-the-art solutions, we have selected 34 existing bot detection methods, each utilizing a diverse range of features. Our comparative evaluation of BotArtist and these existing methods, conducted across nine public datasets under standardized conditions, reveals that the proposed model outperforms existing solutions by almost 10% in terms of F1-score achieving an average score of 83.19 and 68.5 over specific and general approaches respectively. Following the acceptance of this paper, we will make both the developed SAMLP pipeline and the BotArtist model publicly accessible.
Introduction
@@
Online social media has become an essential part of everyday life. During the past decade, online social platforms have managed to transform the communication routine of our daily lives. Due to their growing popularity, online social media gained millions of daily active users that not only consume the information but also create a space for content creators. The main reason behind online social media is real-time access to unlimited information, where registered users can share their comments and personal opinions about popular topics. Such high interest in online social phenomena generates the opportunity and the need for different categories of human interactions and content-sharing platforms. These opportunities lead to the creation of different online social platforms each of which provides a unique user experience, with a very similar goal of real-time human communication and content sharing. Twitter, one of the most popular social networks, with millions of active users, is used for news dissemination, political discussions, and social interactions. However, the platform has also been plagued by the presence of bots and fake accounts used to manipulate and spread false information. According to the research community, the usage of manipulation techniques implemented with the use of bot accounts is registered during diverse popular topic discussions. More specifically, studies show that bot accounts are involved in discussions around the 2016 and 2020 US Presidential elections (Golovchenko et al. 2020; Badawy, Ferrara, and Lerman 2018; Howard, Kollanyi, and Woolley 2016; Shevtsov et al. 2022, 2023b). Besides the US elections a high bot activity with spreading of misleading information is also detected during election periods (presidential/parliamentary/state) in countries like Germany, Sweden, France, Spain, etc. (Neudert, Kollanyi, and Howard 2017; Pastor-Galindo et al. 2020; Bradshaw et al. 2017; Fernquist, Kaati, and Schroeder 2018; Castillo et al. 2019; Rossi et al. 2020). Furthermore, election discourse is not the only point of interest for bot account creators. Spreading propaganda, advertisement, and fake news are identified during the vaccination debate (Broniatowski et al. 2018), the advertisement of e-cigarette (Allem and Ferrara 2016; Allem et al. 2017) and more recent examples of COVID-19 pandemic (Shahi, Dirkson, and Majchrzak 2021; Ferrara 2020; Yang et al. 2021). This high activity of bot accounts raises concerns in the research community and online social media platforms about the integrity of the shared information. Twitter has been taking measures to detect and suspend these bots to maintain the integrity of the platform. Most of the published studies utilize advanced ML techniques to analyze the content and features (structural/network).
@@
Currently, the research community offers a variety of machine learning and neural network-based bot detection methods. Each of these provided methods excels at addressing specific bot detection scenarios, yielding optimal solutions for particular use cases. Unfortunately, the performance of these provided methods significantly diminishes in more general bot detection scenarios, encompassing different periods, discussion topics, and languages, among others. Recent studies have demonstrated that existing methods still fall short of achieving flawless Twitter bot detection, regardless of the number of user characteristics they incorporate.
@@
In our research, we address this challenge by developing a semi-automatic machine learning pipeline (SAMLP) for constructing a generic bot detection model. Next, we compare this developed model with 34 currently available state-of-the-art approaches using nine publicly available datasets. This research yields two distinct evaluation scenarios: one specific to each dataset, where each model is trained and tested on each dataset separately, and a more generic approach where the models are trained and tested on the combined dataset comprising all nine unique datasets.
@@
As a result of our research, we introduce BotArtist, a finely tuned general bot detection model that achieves the highest F1-score performance on three out of the nine datasets, with an average F1-score performance of 83.19. Additionally, the presented approach achieves greater accuracy in the general evaluation, with an overall F1-score improvement of more than 9% and an average improvement of 5% compared to the best-performing existing bot detection approaches. Moreover, BotArtist relies on only a limited set of profile features, allowing for the prediction of historical data independently of the Twitter API.
Finally, we offer insights into the decision-making process of BotArtist through the use of the SHAP method. This approach enables us to explain the model’s decisions and highlight the distinctions between normal and bot accounts as identified by the developed model. Upon paper acceptance, we will make the SAMLP pipeline and the generated BotArtist model publicly accessible.
Related Work
@@
Methods falling into this category perform feature engineering based on information extracted from Twitter user profiles and their activity patterns. Researchers utilize traditional machine learning or neural network classification algorithms for bot detection. These approaches employ different sets of user characteristics as feature sets, including user metadata (Kudugunta and Ferrara 2018), tweets (Miller et al. 2014), description (Hayawi et al. 2022), temporal patterns (Mazza et al. 2019), and follow relationships (Feng et al. 2021a). Moreover, some aim to enhance the scalability of feature-based methods (Yang et al. 2020), discover unknown bot accounts using correlation-based techniques (Chavoshi, Hamooni, and Mueen 2016), and improve the trade-off between precision and recall in bot detection (Morstatter et al. 2016). However, creators of bot accounts are increasingly aware of the features utilized by the research community, leading to novel bot implementations designed to evade detection based on known features (Cresci 2020). Consequently, existing feature-based methods face several challenges in the accurate detection of these new bot accounts (Feng et al. 2021a).
Text based
@@
Text-based methods rely on natural language processing (NLP) techniques for Twitter bot detection, extracting characteristics from posted content (tweets) and user profile descriptions. Approaches in this category include word embeddings (Wei and Nguyen 2019), recurrent neural networks (RNN) (Kudugunta and Ferrara 2018), attention mechanisms (Feng et al. 2021a), and the adoption of pre-trained language models for encoding tweets (Dukić, Keča, and Stipić 2020). Researchers have also combined tweet representations with user profile characteristics (Cai, Li, and Zengi 2017), employed unsupervised machine learning models (Feng et al. 2021a), and addressed multilingual context issues (Knauth 2019). Despite the extensive existing research studies and impressive performances in text-based approaches, novel bot accounts can still evade detection by sharing stolen content from genuine users (Cresci 2020). Additionally, recent work has demonstrated that relying solely on textual information is insufficient for robust and accurate bot detection (Feng et al. 2021c).
Graph based
@@
The graph-based category of bot detection methods combines geometric deep neural networks with graph analytics. Current implementations leverage techniques such as node centrality (Dehghan et al. 2023), node representation learning (Pham et al. 2022), graph neural networks (GNNs) (Ali Alhosseini et al. 2019), and heterogeneous GNNs (Feng et al. 2021c) to conduct graph-based bot detection. In recent research, authors have borrowed ideas from other categories to merge different approaches, such as combining graph and text-based methods (Guo et al. 2021b; Yang, Ferrara, and Menczer 2022). Furthermore, novel GNN architectures have been proposed to exploit the heterogeneity in the Twitter network (Feng et al. 2022a). Generally, these approaches hold significant promise for Twitter bot detection.
@@
Unfortunately, existing Twitter bot detection methods have their limitations. Simplified feature-based approaches struggle to generalize and provide robust results, while more complex methods based on text or graph characteristics require intensive computational resources and large datasets for model development. These limitations, coupled with the recent announcement of the monetizing Twitter API (Twitter 2023), consist of many existing methods expensive to maintain or nearly impossible to operate daily. In this research, we address these challenges and present robust and accurate Twitter bot detection methods that rely on a lightweight set of features. Our methods require only a single user profile object request and can predict users using historical user objects (Twitter API v1.1 or v2) without additional Twitter API requests, thereby significantly reducing operational costs.
Datasets
@@
For this research paper, we have collected nine well-known publicly available datasets (Feng et al. 2021b). All the selected datasets already contain ground truth labels, primarily obtained through manual analysis or crowd-sourcing. To simplify, we have labeled the selected datasets as follows: C-15 (Cresci et al. 2015), G-17 (Gilani et al. 2017), C-17 (Cresci et al. 2017b, a), M-18 (Yang et al. 2020), C-S-18 (Cresci et al. 2018, 2019), C-R-19 (Mazza et al. 2019), B-F-19 (Yang et al. 2019), TwiBot-20 (Feng et al. 2021b), and TwiBot-22 (Feng et al. 2022b). In table: 1, we present the information provided in each dataset, along with the volume of normal and bot accounts.
Dataset C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TwiBot-20 TwiBot-22
# Total User 5,301 2,484 14,368 50,538 13,276 693 518 229,580 1,000,000
# Human 1,950 1,394 3,474 8,092 6,174 340 380 5,237 860,057
# Bot 3,351 1,090 10,894 42,446 7,102 353 138 6,589 139,943
# Total Tweet 2,827,757 0 6,637,615 0 0 0 0 33,488,192 88,217,457
# Human Tweet 2,631,730 0 2,839,361 0 0 0 0 33,488,192 81,250,102
# Bot Tweet 196,027 0 3,798,254 0 0 0 0 33,488,192 6,967,355
# Graph Edges 7,086,134 0 6,637,615 0 0 0 0 33,488,192 170,185,937
Table 1: Description of the selected datasets and the information contained in those datasets.
@@
Figure 1: Machine learning pipeline used for BotArtist model creation. Includes feature selection and model fine-tuning; each step is executed during separate K-Fold cross-validation.
Methodology
@@
The proposed approach is based on the development of a semi-automatic machine learning pipeline (SAMLP). We have chosen this implementation due to its simplicity for further usage and its ability to avoid many trivial mistakes during data processing, feature selection, hyper-parameter fine-tuning, and model evaluation. The steps of the developed pipeline are presented in figure: 1, where the initial step involves data splitting.
@@
During data splitting, it is crucial to maintain a class balance between the training/validation and testing data portions. For this purpose, we utilize a stratified data split with a ratio of 70% for training/validation and 30% for the testing data portions, respectively. This data separation allows us to keep exactly 70% of each class during the training, while preserving the natural class distribution differences, present in the dataset. The testing data portion is kept hidden and is only utilized during the final testing of the model.
Feature Selection
@@
After the data splitting, our pipeline applies a feature selection procedure over the train/validation data portion. For this purpose, we use the K-Fold cross-validation approach, which allows us to utilize the same data portion for multiple trainings and evaluations. Although K-Fold cross-validation may lead to slightly over-optimistic performance estimations, in our case, it is not relevant since we use this procedure for alpha hyper-parameter fine-tuning of the Lasso model, and we are interested in identifying the best alpha parameter.
@@
Additionally, since the Lasso regression model does not perform well over imbalanced datasets, we also take into account the class imbalance of the train/validation dataset. When the class balance is not equal, we utilize under-sampling of the majority class, making the data perfectly balanced and improving the prediction performance of the Lasso model. Due to the high-class imbalance and the under-sampling of the majority class, there is a very high probability of losing some important samples, which may lead to inaccurate feature selection. To address this, we utilize 10 executive repetitions to capture as many samples of the majority class as possible.
@@
During these repetition rounds, we store the best alpha parameters based on the square mean error metric from the entire stratified k-Fold cross-validation results, where $K=5$. At the end of the procedure, we compute the most frequent best $\alpha$ parameter and train the Lasso model with the entire train/validation dataset, keeping features with non-zero coefficients. Additionally, we check if the selected $\alpha$ parameter falls within the defined limits of the searching area. If it falls on the minimum or maximum values, we create a new searching area based on the original site to ensure the selected $\alpha$ is genuinely the best.
@@
This approach allows us to utilize feature selection without additional knowledge of the original data and modifications, as the procedure automatically finds the best alpha values, manages class balance and imbalance via under-sampling and repeated executions, and processes both perfectly balanced and highly imbalanced datasets without information loss.
Model Fine-Tuning

### v4

@@ Abstract
Twitter, as one of the most popular social networks, provides a platform for communication and online discourse. Unfortunately, it has also become a target for bots and fake accounts, resulting in the spread of false information and manipulation. This paper introduces a semi-automatic machine learning pipeline (SAMLP) designed to address the challenges correlated with machine learning model development. Through this pipeline, we develop a comprehensive bot detection model named BotArtist, based on user profile features. SAMLP leverages nine distinct publicly available datasets to train the BotArtist model. To assess BotArtist’s performance against current state-of-the-art solutions, we select 35 existing Twitter bot detection methods, each utilizing a diverse range of features. Our comparative evaluation of BotArtist and these existing methods, conducted across nine public datasets under standardized conditions, reveals that the proposed model outperforms existing solutions by almost 10%, in terms of F1-score, achieving an average score of 83.19 and 68.5 over specific and general approaches respectively. As a result of this research, we provide a dataset of the extracted features combined with BotArtist predictions over the 10.929.533 Twitter user profiles, collected via Twitter API during the 2022 Russo-Ukrainian War, over a 16-month period. This dataset was created in collaboration with (Shevtsov et al., 2022a) where the original authors share anonymized tweets on the discussion of the Russo-Ukrainian war with a total amount of 127.275.386 tweets. The combination of the existing text dataset and the provided labeled bot and human profiles will allow for the future development of a more advanced bot detection large language model in the post-Twitter API era.
1 Introduction
Online social media has become an essential part of everyday life. During the past decade, social networks have transformed the communication routine of our daily lives. Due to their growing popularity, online social media gained millions of daily active users not only consuming the information but also creating a space for content creators. The main reason behind the success of online social networks is the real-time access to unlimited information, where registered users can share their comments and personal opinions on popular topics.
@@ 1 Introduction
Twitter, one of the most popular social networks, with millions of active users, is used for news dissemination, political discussions, and social interactions. However, the platform has also been plagued by bots and fake accounts that are used to manipulate and spread false information. According to the research community, the usage of manipulation techniques implemented with the use of bot accounts is registered during diverse popular topic discussions. More specifically, studies show that bot accounts are involved in diverse political discussions, for instance, around 2016 in countries like the US, Germany, Sweden, France, Spain, etc. (Golovchenko et al., 2020; Badawy et al., 2018; Howard et al., 2016; Shevtsov et al., 2022b, 2023; Neudert et al., 2017; Pastor-Galindo et al., 2020; Bradshaw et al., 2017; Fernquist et al., 2018; Castillo et al., 2019; Rossi et al., 2020). Furthermore, propagandizing and false news is identified during the vaccination debate (Broniatowski et al., 2018) and more recent examples of the COVID-19 pandemic (Shahi et al., 2021; Ferrara, 2020; Yang et al., 2021). This high activity of bot accounts raises concerns in the research community and online social media platforms about the integrity of the shared information.
@@
Currently, the research community offers a variety of ML and NN-based bot detection methods. Each of these provided methods excels at addressing specific bot detection scenarios, yielding optimal solutions for particular use cases. Unfortunately, the performance of these provided methods significantly diminishes in more general bot detection scenarios, encompassing different periods, discussion topics, and languages, among others. Recent studies have shown that existing methods still do not achieve flawless Twitter bot detection, regardless of the number of user characteristics that they incorporate (Feng et al., 2022a). In addition, a large portion of existing ML bot detection methods ignore well-known optimization procedures (including feature selection and dynamic hyper-parameter fine-tuning) providing significant space for improvement.
@@
In our research, we address this challenge by developing a semi-automatic machine learning pipeline (SAMLP [^1] GitHub repository: TBD ) for constructing a generic bot detection model (BotArtist [^2] GitHub repository: TBD ). The presented pipeline resolves popular issues during the creation of the classification model, including recursive hyperparameter fine-tuning during the feature selection phase. This approach guarantees the reduction of non-relevant features which sequentially reduces the noise of initial data. Additionally, our pipeline considers the class imbalance during feature selection and model fine-tuning, making the model suitable for real-case applications. Next, we compare this developed model with 35 currently available state-of-the-art approaches using nine publicly available datasets. This research yields two distinct evaluation scenarios: one specific to each dataset, where each model is trained and tested separately on each of them, and a more generic approach, where the models are trained and tested on the combined dataset comprising all nine unique datasets.
@@
As a result of our research, we introduce BotArtist, a finely tuned general bot detection model that achieves the highest F1-score performance on three of the nine datasets, with an average F1-score performance of 83.19. Furthermore, the presented approach achieves greater precision in general evaluation, with an overall improvement in F1 score of more than 9% and an average improvement of 5% compared to the best-performing existing bot detection approaches. Moreover, BotArtist relies on only a limited set of profile features, allowing for the prediction of historical data independently of the Twitter API.
Except from the source code of SAMLP and BotArtist, we make available the model outcome usable for research. For this purpose, we collect a large volume of users correlated to the already existing textual shared dataset and add an additional layer to textual information via user prediction. We then release 10.929.533 user profiles with BotArtist predictions  [^3] Zenodo repository: TBD , correlated to 127.275.386 tweets (Shevtsov et al., 2022a).
2 Related Work
@@ 2.1 Feature based
Methods falling into this category apply feature engineering based on information extracted from Twitter user profiles and their activity patterns. Researchers use traditional machine learning or neural network classification algorithms for bot detection. These approaches employ different sets of user characteristics as feature sets, including user metadata (Kudugunta and Ferrara, 2018), tweets (Miller et al., 2014), user name (Beskow and Carley, 2019), description (Hayawi et al., 2022), temporal patterns (Mazza et al., 2019), and follow relationships (Feng et al., 2021a). Moreover, some of them aim to enhance the scalability of feature-based methods (Yang et al., 2020; Abreu et al., 2020), discover unknown bot accounts using correlation-based techniques (Chavoshi et al., 2016), and improve the trade-off between precision and recall in bot detection (Morstatter et al., 2016). However, the creators of bot accounts are increasingly aware of the features utilized by the research community, leading to novel bot implementations designed to evade detection based on known features (Cresci, 2020). Consequently, existing feature-based methods face several challenges in the accurate detection of these new bot accounts (Feng et al., 2021a).
2.2 Text based
@@
Text-based methods rely on natural language processing (NLP) techniques for Twitter bot detection, extracting characteristics from posted content (tweets) and user profile descriptions. Approaches in this category include sequence fingerprint (Cresci et al., 2016), word embeddings (Wei and Nguyen, 2019), recurrent neural networks (RNN) (Kudugunta and Ferrara, 2018), attention mechanisms (Feng et al., 2021a), transformers (Guo et al., 2021a; Liu et al., 2019; Raffel et al., 2020) and the adoption of pre-trained language models for encoding tweets (Dukić et al., 2020). Researchers have also combined tweet representations with user profile characteristics (Cai et al., 2017; Efthimion et al., 2018; Kantepe and Ganiz, 2017; Miller et al., 2014; Varol et al., 2017; Kouvela et al., 2020; Lee et al., 2011; Echeverria et al., 2018), employed unsupervised machine learning models (Feng et al., 2021a), and addressed multilingual context issues (Knauth, 2019). Despite extensive existing research studies and impressive performance in text-based approaches, new bot accounts can still evade detection by sharing stolen content from genuine users (Cresci, 2020). Additionally, recent work has demonstrated that relying solely on textual information is insufficient for robust and accurate bot detection (Feng et al., 2021b).
2.3 Graph based
@@
The graph-based category of bot detection methods combines geometric deep neural networks with graph analytics. Current implementations leverage techniques such as node centrality (Dehghan et al., 2023), node representation learning (Pham et al., 2022), graph neural networks (GNNs) (Ali Alhosseini et al., 2019; Moghaddam and Abbaspour, 2022), and heterogeneous GNNs (Feng et al., 2021b) to conduct graph-based bot detection. In recent research, the authors have borrowed ideas from other categories to merge different approaches, such as combining multiple methods (Guo et al., 2021b; Yang et al., 2022; Knauth, 2019; Beskow and Carley, 2020; Rodríguez-Ruiz et al., 2020; Magelinski et al., 2020; Yang et al., 2013; Kipf and Welling, 2016, 2016; Veličković et al., 2017; Lv et al., 2021). Furthermore, novel GNN architectures have been proposed to exploit heterogeneity in the Twitter network (Feng et al., 2022b). Generally, these approaches hold significant promise for Twitter bot detection.
@@
Unfortunately, existing Twitter bot detection methods have their limitations. Simplified feature-based approaches struggle to generalize and provide robust results, while more complex methods based on text or graph characteristics require intensive computational resources and large datasets for model development. Furthermore, as shown in (Shevtsov et al., 2024) the complex embedding-based model tends to perfectly capture the training data patterns, but lacks generalizability and requires frequent retraining over multiple datasets to keep high performance. These limitations, coupled with the recent announcement of the Twitter API monetization (Twitter, 2023), make many existing methods expensive to maintain or nearly impossible to operate daily. In this study, we address these challenges and present robust and accurate Twitter bot detection methods, based on a lightweight set of features. Our methods require only a single user profile object (Twitter API v1.1 or v2) to provide user prediction without additional Twitter API requests, thereby significantly reducing operational costs.
3 Datasets
@@
For this research paper, we collect nine well-known publicly available datasets (Feng et al., 2021c). All selected datasets already contain ground truth labels, primarily obtained through manual analysis or crowd-sourcing. For simplicity, we label the selected datasets as follows: C-15 (Cresci et al., 2015), G-17 (Gilani et al., 2017), C-17 (Cresci et al., 2017a, b), M-18 (Yang et al., 2020), C-S-18 (Cresci et al., 2018, 2019), C-R-19 (Mazza et al., 2019), B-F-19 (Yang et al., 2019), TwiBot-20 (Feng et al., 2021c), and TwiBot-22 (Feng et al., 2022a). In Table 1, we present the information provided in each dataset, along with the volume of normal and bot accounts.
Furthermore, in collaboration with (Shevtsov et al., 2022a) collect 10.929.533 Twitter profiles correlated with 127.275.386 publicly available tweets related to the public discussion topic of the 2022 Russo-Ukrainian War. Our collection of user profiles is based on the monitoring of selected topics starting from February 23, 2022, till June 23, 2023. The shared datasets contain a set of extracted features in an anonymized form of a CSV file and contain only preprocessed numerical features to protect user information. The provided dataset also provides anonymized user IDs which are identically correlated with publicly available user tweet dataset (Shevtsov et al., 2022a).
Table 1: Description of selected datasets and information contained in those datasets.
Dataset C-15 G-17 C-17 M-18 C-S-18 C-R-19 B-F-19 TwiBot-20 TwiBot-22 # Total User 5,301 2,484 14,368 50,538 13,276 693 518 229,580 1,$0$ Human 1,950 1,394 3,474 8,092 6,174 340 380 5,237 860,057 # Bot 3,351 1,090 10,894 42,446 7,102 353 138 6,589 139,943 # Total Tweet 2,827,757 0 6,637,615 0 0 0 0 33,488,192 88,217,457 # Human Tweet 2,631,730 0 2,839,361 0 0 0 0 927,292 81,250,102 # Bot Tweet 196,027 0 3,798,254 0 0 0 0 1,072,496 6,967,355 # Graph Edges 7,086,134 0 6,637,615 0 0 0 0 33,488,192 170,185,937
@@
Figure 1: Machine learning pipeline used for BotArtist model creation including feature selection and model fine-tuning; each step is executed during separate K-Fold cross-validation.
4 Methodology
@@
The proposed approach is based on the development of a semi-automatic machine learning pipeline (SAMLP). We choose this implementation due to its simplicity for further usage and its ability to avoid many trivial mistakes during the ML model development such as data processing, feature selection, hyper-parameter fine-tuning, and model evaluation. The steps of the developed pipeline are presented in Figure 1, where the initial step involves data splitting.
@@
During data splitting, it is crucial to maintain a class balance between the training/validation and testing data portions. For this purpose, we apply a stratified data split with a 70:30 ratio for training, validation, and testing portions respectively. Such an approach allows us to preserve the natural class distribution difference. The testing data portion is kept hidden and is only utilized during the final testing of the model to prevent information leakage between train and test portions
4.1 Feature Selection
@@
After the data splitting, our pipeline applies a feature selection procedure over the train/validation data portion during the K-Fold cross validation. Although K-Fold cross-validation may lead to slightly over-optimistic performance estimations, in our case, it is not relevant since we use this procedure for $\alpha$ hyper-parameter fine-tuning of the Lasso model, and we are interested in identifying the best $\alpha$ parameter.
@@
Additionally, since the Lasso regression model does not perform well over imbalanced datasets, we also take into account the class imbalance of the dataset. In case of data imbalance,we utilize under-sampling of the majority class, making the data perfectly balanced and improving the prediction performance of the Lasso model. Due to the high-class imbalance and the under-sampling of the majority class, there is a very high probability of losing some important samples, which may lead to inaccurate feature selection. To address this, we utilize 10 consecutive repetitions to capture as many samples of the majority class as possible.
@@
During these repetition rounds, we store the best $\alpha$ parameters based on the square mean error metric from the entire stratified k-Fold cross-validation results, where $K=5$. At the end of the procedure, we compute the most frequent best $\alpha$ parameter and train the Lasso model with the entire train/validation dataset, keeping features with non-zero coefficients. Additionally, we check if the selected $\alpha$ parameter falls within the defined limits of the searching area. If it falls on the minimum or maximum values, we create a new search area based on the original site to ensure the selected $\alpha$ is genuinely the best.
@@
This approach allows us to utilize feature selection without additional knowledge of the original data and modifications, as the procedure automatically finds the best $\alpha$ values, manages class balance and imbalance via under-sampling and repeated executions, and processes both perfectly balanced and highly imbalanced datasets without information loss.
4.2 Model Fine-Tuning